# Learning Unified User Quantized Tokenizers for User Representation

### v1

@@ Abstract
Multi-source user representation learning plays a critical role in enabling personalized services on web platforms (e.g., Alipay). While prior works have adopted late-fusion strategies to combine heterogeneous data sources, they suffer from three key limitations: lack of unified representation frameworks, scalability and storage issues in data compression, and inflexible cross-task generalization. To address these challenges, we propose U²QT (Unified User Quantized Tokenizers), a novel framework that integrates cross-domain knowledge transfer with early fusion of heterogeneous domains. Our framework employs a two-stage architecture: first, a causal Q-Former projects domain-specific features into a shared causal representation space to preserve inter-modality dependencies; second, a multi-view RQ-VAE discretizes causal embeddings into compact tokens through shared and source-specific codebooks, enabling efficient storage while maintaining semantic coherence. Experimental results showcase U²QT’s advantages across diverse downstream tasks, outperforming task-specific baselines in future behavior prediction and recommendation tasks while achieving efficiency gains in storage and computation. The unified tokenization framework enables seamless integration with language models and supports industrial-scale applications.
Introduction
@@
In the web platforms such as Alipay111https://global.alipay.com/platform/site/ihome, users generate diverse multi-source information, including textual profiles, behavioral histories, and structured tabular attributes. Effectively leveraging this heterogeneous data is crucial for advancing personalized services and optimizing user experience. To this end, prior studies have predominantly adopted a late-fusion strategy, wherein source-specific features are independently extracted and subsequently combined at a later stage (Purificato, Boratto, and De Luca 2024; Wu et al. 2019; Tang et al. 2021). For example, early approaches, such as user2Vec (Hallac et al. 2019), adapted document embedding models (e.g., doc2vec) to encode social media users, typically by aggregating user vectors within the same category to form category-level representations. To address the limitation of simple aggregation, subsequent works designed more expressive fusion mechanisms. For instance, MFN (Bian et al. 2021) introduced a Macro-micro Fusion Network that captures semantic alignment at the user level using a category-based fusion component. In parallel, attention-based models like MSSR (Lin et al. 2024) explored how to jointly leverage both inter- and intra-sequence user interactions to capture a user’s diverse behavioral patterns.
@@
With the rapid growth of large-scale pre-training, general-purpose user representation frameworks have emerged. For example, One4all (Shin et al. 2021) learns universal user embeddings via large-scale pre-training, and its extension U-MLP further enhances user targeting via an MLP-based decoder. Sequence modeling methods such as MSDP (Fu et al. 2023) focus on modeling temporal dynamics of user behaviors at multiple scales.
@@
Recently, to support zero-shot transfer and few-shot user targeting, FOUND (Dou et al. 2025) proposed a self-supervised pre-training framework with user-text alignment, advancing user modeling towards more generalizable and transferable representations.
@@
(i) Lack of unified representation for heterogeneous data: Methods like User2Vec overlook cross-source interactions by relying on single-source features, while existing approaches (e.g., MSSR, MFN) employ late fusion strategies in multi-source integration, limiting cross-source interaction potential.
(ii) Data utilization and storage issues: Methods like FOUND require massive scale to store embeddings for every user, leading to prohibitive storage demands and further impede the model’s capacity to fully capitalize on abundant information.
(iii) Inflexible cross-task generalization: Methods such as Conure (Yuan et al. 2021) and TERACON (Kim et al. 2023) require additional fine-tuning of a large number of user embeddings when adapted to new tasks, resulting in limited flexibility and restricted practical applicability.
To address the above challenges, we propose Unified User Quantized Tokenizers (U2QT), which compresses the heterogeneous multi-source data into a unified quantitative language for efficient data storage, utilization and computation. Specifically, we adopt a two-stage quantitative language-based data compression framework to solve Challenge (i) and (ii). In the first stage, heterogeneous source-specific features (e.g., text, sequential behaviors, tabular data) are projected into a shared transformer encoder. A causal Q-Former then maps these features into unlearned query vectors, effectively transforming them into a unified causal representation space that preserves inter-source dependencies. In the second stage, we propose a MRQ-VAE (Multi-View Residual Quantized Variational AutoEncoder) to discretize the causal embeddings into compact tokens stored in shared and source-specific codebooks where cross-source semantic coherence and source-unique patterns can be learned. This quantization process achieves 84× memory reduction compared to FOUND (Dou et al. 2025), with 3.5× faster training speed under equivalent training data scales, while enhancing data utilization efficiency through semantic-preserving token compression. The method enables scalable deployment of large-scale user behavior data without compromising semantic fidelity. Furthermore, due to its lightweight and quantized design, which eliminates redundant and irrelevant information, our method achieves strong cross-task adaptability. To assess the effectiveness and cross-task adaptability of our approach, we introduce a multi-task e-commerce benchmark encompassing user profile, future behavior prediction, and downstream recommendation tasks. Extensive experiments demonstrate superior adaptability and scalability across diverse tasks compared to state-of-the-art baselines. In summary, our contributions are listed as follows,
@@
Related Work
Quantitative Language
Quantitative language has been widely used in VLLM (Li et al. 2025), generative recommendation and retrieval (Li et al. 2024), which is also called semantic IDs (Zheng et al. 2024; Pan et al. 2024; Zhang et al. 2025). For instance, VQ-Rec (Hou et al. 2023) proposes to generate code using content information for item representation. Tiger (Rajput et al. 2023) utilizes RQ-VAE to generate semantic IDs for hierarchical item representation. To represent multi-modal features, MMGRec (Liu et al. 2024) utilize graph RQ-VAE to generate semantic IDs for multi-modal item representation fused by GCN. MQL4GRec (Zhai et al. 2025) transforms item content from different domains and modalities into a unified quantitative language to facilitate the effective transfer of recommendation knowledge. SEED (Ge et al. 2023) proposes a discrete image tokenizer, designed based on the premise that visual tokens compatible with LLMs should capture high-level semantics while being generated with a 1D causal dependency.
Universal User Representation
User modeling, also known as user profiling, refers to the process of collecting, extracting, and representing user features and personal characteristics to construct comprehensive and accurate models of users(Purificato, Boratto, and De Luca 2024). Existing works has focused on the creating universal user representation(Yuan et al. 2021; Shin et al. 2021). DUPN (Ni et al. 2018) perform multitask representation learning using an attention-based RNN architecture to capture in-depth representations of portal users. PeterRec (Yuan et al. 2020) propose a parameter-efficient transfer learning framework that reduces computational costs in fine-tuning. It supports five downstream tasks, including user profile prediction and cold-start recommendations for browsers. SUMN (Gu et al. 2021) propose a Self-supervised User Modeling Network to learn universal user representations from behavior data. It consists of a self-supervised learning objective to capture and preserve meaningful user information, along with a multi-hop aggregation layer that enhances the model’s ability to integrate diverse behavioral signals. Conure (Yuan et al. 2021) propose the first continual user representation learning framework that adapts to new tasks while retaining knowledge from previous ones through iterative pruning of less important model weights. TERACON (Kim et al. 2023) propose a continual user representation learning method that captures task relationships via task-specific soft masks. A knowledge retention module with pseudo-labeling helps prevent catastrophic forgetting during sequential learning.
Preliminaries
Notations
@@
where is the multi-source user data used for pretraining, represents the pretrained unified user representation, and denotes the proposed pretraining task, which will be illustrated in the Section Methodology.
@@
where is the predicted label of -th user in the downstream task , is the model parameters for specific tasks. The specific application of our pretrained unified user representation to downstream tasks will be detailed in Section Cross-Task Application. In summary, this is a two-stage framework in which user representations are generated first and fed into a specific model for downstream task.
Methodology
@@
In this section, we elaborate on the proposed Unified User Quantized Tokenizers, a novel approach of user modeling. First, we provide an overview of U2QT in Section Overview of Unified User Tokenizer to facilitate a thorough understanding of our method. Subsequently, we illustrate the detail architecture of our proposed method including Multi-source Encoding (Section Multi-source Encoding), Cross-source Fusion (Section Cross-source Fusion and Condensation), and Quantized User Tokenizers (Section Quantized User Tokenizers). Finally, we show the flexible application of user tokenizers in future behavior prediction and downstream recommendation tasks in Section Cross-Task Application. Comprehensive architecture of U2QT is illustrated in Figure 1.
Overview of Unified User Tokenizer
@@
In this subsection, we illustrate the process of transforming complex mutli-source content of user into a concise and unified quantitative language. To better construct a unified and rich-semantic user tokenizer, we introduce a two-stage compression process, including continuous latent space and discrete embedding space compression. The first stage ensures an effective filter of task-related semantic among multi-source heterogeneous data, and the second stage contributes to efficient storage and advanced data utilization proficiency.
Multi-source Encoding
Considering the heterogeneity and information disparity across different data sources, we process each source independently and encode them into latent representations to streamline subsequent procedures. To enable efficient encoding of multi-source data, we adopt a shared Transformer-based encoder. Formally, we encode preprocessed heterogeneous user data using a shared encoder based on transformer architecture (Huang et al. 2020; Devlin et al. 2019), as follows:
where denotes Transformer Encoder, and represent embeddings of user behavior, tabular data, and search-related information, respectively.
Cross-source Fusion and Condensation
Different source data often share common knowledge that can be mutually beneficial. Effectively leveraging this shared information to enable cross-source complementarity is crucial for enhancing and condensing representations. In this subsection, we propose a cross-source communication mechanism that promotes mutual enrichment among sources while concurrently compressing source-specific representations into a compact latent space. To this end, we employ the Q-Former (Querying Transformer) (Li et al. 2023) to transform multi-source heterogeneous data into unified, learnable query vectors within a compact continuous latent space, thereby enabling efficient semantic condensation and cross-source communication. Formally, we take the source-specific user representations as input and leverage the cross-attention mechanism within the Q-Former to extract and condense task-relevant semantic patterns into learnable query vectors, while simultaneously facilitating cross-source communication:
where denotes the shared learnable query vectors across different sources, facilitating cross-source communication. Thus, the multi-source user representations are compressed into unified continuous latent spaces , where denotes the number of layers in the Q-Former.
Quantized User Tokenizers
@@
The quantized representation is then fed into a reverse Q-Former to reconstruct the continuous latent user representation, followed by a Transformer decoder to reconstruct the raw user content . For notational simplicity, we also omit symbols indicating source information and whether the codebook is shared or specific. The loss function comprises:
@@
Cross-Task Application
@@
Future Behavior Prediction
In this subsection, we aim to endow our user tokenizer with the future behavior prediction capabilities. Inspired by CLIP and it’s variants (Li et al. 2022, 2023; Zhu et al. 2025), we align user tokenizer representations with future behaviors via contrastive learning for prediction. Future behavior descriptions are encoded into semantic embeddings using a LoRA-tuned (Hu et al. 2022) LLM:
@@
After obtaining the fused representation of user tokenizer and corresponding text embedding of future user behavior in natural langugage semantic , we employ the contrastive loss Info-NCE (van den Oord, Li, and Vinyals 2019) to align these two representations in different semantic space. The loss is applied to pulling together the two representations from same user while pushing away two representations from different users. The contrastive loss is performed as follow:
@@
Recommendation Task
In this section, we illustrate the usage of user tokenizer in downstream recommendation task. To describe the user-side feature, we employ the unified user tokenizer as a part of model input to representation multifaceted user information extracted from heterogeneous multi-source data. The overall process is as follows:
where denotes the user tokenizer of extracted from cashe storage, denotes the feature representation of and (Wang, She, and Zhang 2021) is a classical method in CTR prediction.
Experiment
@@
Experimental Setup
Datasets
@@
For pre-training the industrial framework of unified user tokenizers, we employ the real-world industrial dataset, which collects heterogeneous multi-scenario user data on Alipay. Specifically, the training dataset contains around 20 million pieces of user behavior, each piece organized according to Notations. The user representation test benchmark contains 4 scenarios that cover security risk control, future behavior prediction, and the user profile prediction. Specifically, the test dataset contains around 50,000 pieces of user behaviors.
@@
For the downstream recommendation task, we adopt a chronological partitioning strategy: interactions from the first 21 days (nearly 1,550,000 samples) form the training set, while the subsequent 3 days (nearly 100,000 samples) are evenly split into validation and test sets via stratified random sampling. This temporal split strictly avoids data leakage by preserving interaction order and ensures robustness through dedicated hold-out periods.
Baselines

### v2

@@ Abstract.
Multi-source user representation learning plays a critical role in enabling personalized services on web platforms (e.g., Alipay). While prior works have adopted late-fusion strategies to combine heterogeneous data sources, they suffer from three key limitations: lack of unified representation frameworks, scalability and storage issues in data compression, and inflexible cross-task generalization. To address these challenges, we propose U²QT (Unified User Quantized Tokenizers), a novel framework that integrates cross-domain knowledge transfer with early fusion of heterogeneous domains. Our framework employs a two-stage architecture: first, we use the Qwen3 Embedding model to derive a compact yet expressive feature representation; second, a multi-view RQ-VAE discretizes causal embeddings into compact tokens through shared and source-specific codebooks, enabling efficient storage while maintaining semantic coherence. Experimental results showcase U²QT’s advantages across diverse downstream tasks, outperforming task-specific baselines in future behavior prediction and recommendation tasks while achieving efficiency gains in storage and computation. The unified tokenization framework enables seamless integration with language models and supports industrial-scale applications.
1. Introduction
@@
In the web platforms such as Alipay111https://global.alipay.com/platform/site/ihome, users generate diverse multi-source information, including textual profiles, behavioral histories, and structured tabular attributes. Effectively leveraging this heterogeneous data is crucial for advancing personalized services and optimizing user experience. To this end, prior studies have predominantly adopted a late-fusion strategy, wherein source-specific features are independently extracted and subsequently combined at a later stage (Purificato et al., 2024; Wu et al., 2019; Tang et al., 2021). For example, early approaches, such as user2Vec (Hallac et al., 2019), adapted document embedding models (e.g., doc2vec) to encode social media users, typically by aggregating user vectors within the same category to form category-level representations. To address the limitation of simple aggregation, subsequent works designed more expressive fusion mechanisms. For instance, MFN (Bian et al., 2021) introduced a Macro-micro Fusion Network that captures semantic alignment at the user level using a category-based fusion component. In parallel, attention-based models like MSSR (Lin et al., 2024) explored how to jointly leverage both inter- and intra-sequence user interactions to capture a user’s diverse behavioral patterns.
@@
With the rapid growth of large-scale pre-training, general-purpose user representation frameworks have emerged. For example, One4all (Shin et al., 2021) learns universal user embeddings via large-scale pre-training, and its extension U-MLP further enhances user targeting via an MLP-based decoder. Sequence modeling methods such as MSDP (Fu et al., 2023) focus on modeling temporal dynamics of user behaviors at multiple scales.
@@
Recently, to support zero-shot transfer and few-shot user targeting, FOUND (Dou et al., 2025) proposed a self-supervised pre-training framework with user-text alignment, advancing user modeling towards more generalizable and transferable representations.
@@
(i) Lack of unified representation for heterogeneous data: As depicted in Figure 1, methods like User2Vec overlook cross-source interactions by relying on single-source features, while existing approaches (e.g., MSSR, MFN) employ late fusion strategies in multi-source integration, limiting cross-source interaction potential.
(ii) Data utilization and storage issues: As shown in Figure 2, methods like FOUND require massive scale to store embeddings for every user, leading to prohibitive storage demands and further impede the model’s capacity to fully capitalize on abundant information.
(iii) Inflexible cross-task generalization: Methods such as Conure (Yuan et al., 2021) and TERACON (Kim et al., 2023) require additional fine-tuning of a large number of user embeddings when adapted to new tasks, resulting in limited flexibility and restricted practical applicability.
To address the above challenges, we propose Unified User Quantized Tokenizers (U2QT), which compresses the heterogeneous multi-source data into a unified quantitative language for efficient data storage, utilization and computation. Specifically, we adopt a two-stage quantitative language-based data compression framework to solve Challenge (i) and (ii). In the first stage, heterogeneous source-specific features (e.g., text, sequential behaviors, tabular data) are projected into a unified language space in Qwen3 embedding model (Zhang et al., 2025b), which transforms these features in the format of long context into compact yet expressive embeddings. In the second stage, we propose a MRQ-VAE (Multi-View Residual Quantized Variational AutoEncoder) to discretize the causal embeddings into compact tokens stored in shared and source-specific codebooks where cross-source semantic coherence and source-unique patterns can be learned. This quantization process achieves 84× memory reduction compared to FOUND (Dou et al., 2025), with 3.5× faster training speed under equivalent training data scales, while enhancing data utilization efficiency through semantic-preserving token compression. The method enables scalable deployment of large-scale user behavior data without compromising semantic fidelity. Furthermore, due to its lightweight and quantized design, which eliminates redundant and irrelevant information, our method achieves strong cross-task adaptability. To assess the effectiveness and cross-task adaptability of our approach, we introduce a multi-task e-commerce benchmark encompassing user profile, future behavior prediction, and downstream recommendation tasks. Extensive experiments demonstrate superior adaptability and scalability across diverse tasks compared to state-of-the-art baselines. In summary, our contributions are listed as follows,
@@
2. Related Work
2.1. Quantitative Language
Quantitative language has been widely used in VLLM (Li et al., 2025), generative recommendation and retrieval (Li et al., 2024), which is also called semantic IDs (Zheng et al., 2024; Pan et al., 2024; Zhang et al., 2025a). For instance, VQ-Rec (Hou et al., 2023) proposes to generate code using content information for item representation. Tiger (Rajput et al., 2023) utilizes RQ-VAE to generate semantic IDs for hierarchical item representation. To represent multi-modal features, MMGRec (Liu et al., 2024) utilize graph RQ-VAE to generate semantic IDs for multi-modal item representation fused by GCN. MQL4GRec (Zhai et al., 2025) transforms item content from different domains and modalities into a unified quantitative language to facilitate the effective transfer of recommendation knowledge. SEED (Ge et al., 2023) proposes a discrete image tokenizer, designed based on the premise that visual tokens compatible with LLMs should capture high-level semantics while being generated with a 1D causal dependency.
2.2. Universal User Representation
User modeling, also known as user profiling, refers to the process of collecting, extracting, and representing user features and personal characteristics to construct comprehensive and accurate models of users(Purificato et al., 2024). Existing works has focused on the creating universal user representation(Yuan et al., 2021; Shin et al., 2021). DUPN (Ni et al., 2018) perform multitask representation learning using an attention-based RNN architecture to capture in-depth representations of portal users. PeterRec (Yuan et al., 2020) propose a parameter-efficient transfer learning framework that reduces computational costs in fine-tuning. It supports five downstream tasks, including user profile prediction and cold-start recommendations for browsers. SUMN (Gu et al., 2021) propose a Self-supervised User Modeling Network to learn universal user representations from behavior data. It consists of a self-supervised learning objective to capture and preserve meaningful user information, along with a multi-hop aggregation layer that enhances the model’s ability to integrate diverse behavioral signals. Conure (Yuan et al., 2021) propose the first continual user representation learning framework that adapts to new tasks while retaining knowledge from previous ones through iterative pruning of less important model weights. TERACON (Kim et al., 2023) propose a continual user representation learning method that captures task relationships via task-specific soft masks. A knowledge retention module with pseudo-labeling helps prevent catastrophic forgetting during sequential learning.
3. Preliminaries
3.1. Notations
@@ 3.2. Problem Definition
where is the multi-source user data used for pretraining, represents the pretrained unified user representation, and denotes the proposed pretraining task, which will be illustrated in the Section 4.
@@
where is the predicted label of -th user in the downstream task , is the model parameters for specific tasks. The specific application of our pretrained unified user representation to downstream tasks will be detailed in Section 4.4. In summary, this is a two-stage framework in which user representations are generated first and fed into a specific model for downstream task.
4. Methodology
@@
In this section, we elaborate on the proposed Unified User Quantized Tokenizers, a novel approach of user modeling. First, we provide an overview of U2QT in Section 4.1 to facilitate a thorough understanding of our method. Subsequently, we illustrate the detail architecture of our proposed method including Multi-source Encoding (Section 4.2), and Quantized User Tokenizers (Section 4.3). Finally, we show the flexible application of user tokenizers in future behavior prediction and downstream recommendation tasks in Section 4.4. Comprehensive architecture of U2QT is illustrated in Figure 3.
4.1. Overview of Unified User Tokenizer
@@
In this subsection, we illustrate the process of transforming complex mutli-source content of user into a concise and unified quantitative language. To better construct a unified and rich-semantic user tokenizer, we introduce a two-stage compression process, including expressive language space and discrete embedding space compression. The first stage ensures a compact yet expressive semantic representation among multi-source heterogeneous data, and the second stage contributes to efficient storage and advanced data utilization proficiency.
4.2. Multi-source Encoding
Given the heterogeneity and information disparity across data sources, prior methods typically processed each source independently and encoded it into a latent representation to streamline downstream procedures (Huang et al., 2020; Devlin et al., 2019). While this design simplifies the pipeline, it leaves a semantic gap across sources that these models struggle to bridge. To address this, we leverage LLMs to map multi-source data into a shared language space, producing richer and better-aligned embeddings. We use the Qwen3 Embedding model (Zhang et al., 2025b) to encode multi-source data. We construct a prompt that combines the query with the corresponding source data and feed it to the model. Specifically, we show an example of SPM data as illustrated in Appendix D.
Formally, final embeddings of multi-source data are derived from the hidden state of the last layer corresponding to [EOS] token as follows:
Where denotes the final representation of the multi-source input, and indexes the source. This representation compresses very long contextual information into a single embedding.
4.3. Quantized User Tokenizers
@@
The quantized representation is then fed into a source-specific MLP decoder to reconstruct the Qwen3 embedding . For notational simplicity, we also omit symbols indicating source information and whether the codebook is shared or specific. The loss function comprises:
@@
4.4. Cross-Task Application
@@
4.4.1. Future Behavior Prediction
In this subsection, we aim to endow our user tokenizer with the future behavior prediction capabilities. Inspired by CLIP and it’s variants (Li et al., 2022, 2023; Zhu et al., 2025), we align user tokenizer representations with future behaviors via contrastive learning for prediction. Future behavior descriptions are encoded into semantic embeddings using a LoRA-tuned (Hu et al., 2022) LLM:
@@ 4.4.1. Future Behavior Prediction
After obtaining the fused representation of user tokenizer and corresponding text embedding of future user behavior in natural langugage semantic , we employ the contrastive loss Info-NCE (van den Oord et al., 2019) to align these two representations in different semantic space. The loss is applied to pulling together the two representations from same user while pushing away two representations from different users. The contrastive loss is performed as follow:
@@
4.4.2. Recommendation Task
In this section, we illustrate the usage of user tokenizer in downstream recommendation task. To describe the user-side feature, we employ the unified user tokenizer as a part of model input to representation multifaceted user information extracted from heterogeneous multi-source data as depicted in Figure 7. The overall process is as follows:
where denotes the user tokenizer of extracted from cashe storage, denotes the feature representation of and (Wang et al., 2021) is a classical method in CTR prediction.
5. Experiment
@@
5.1. Experimental Setup
5.1.1. Datasets
@@
We pre-train our industrial framework of unified user tokenizers on a real-world dataset from Alipay that aggregates heterogeneous, multi-scenario user data. The training set comprises approximately 20,000,000 user-behavior records, each organized according to the Notations. The user-representation benchmark covers four scenarios spanning security risk control, future behavior prediction, and user profile prediction. The test set contains approximately 50,000 user-behavior records.
@@
For the downstream recommendation task, we adopt a chronological partition: interactions from the first 21 days (approximately 1,550,000 samples) form the training set, while the subsequent 3 days (approximately 100,000 samples) are split evenly into validation and test sets using stratified random sampling. This temporal split preserves interaction order to prevent data leakage and enhances robustness through dedicated holdout periods.
5.1.2. Detailed Experimental Setup
The detailed experimental setup, including baseline descriptions, evaluation metrics, and implementation details, is provided in Appendix A, Appendix B, and Appendix C, respectively.
| Feature | Backbone | AUC | HR@10 | HR@20 |
| User Profile | Masknet | 0.6476 | 0.2308 | 0.2968 |
| (+)Tokenizer | Masknet | 0.6527 | 0.2678 | 0.4913 |
5.2. Performance Analysis
As illustrated in Table 2, the experimental results demonstrate that U2QT significantly outperforms baseline methods across downstream tasks in our designed benchmark. Our model achieves superior performance in AUC and KS in all domains and outperforms the second-best baseline 3% / 9% on average in AUC / KS. Moreover, we utilize the pretrained user tokenizer as supplementary feature in existing methods(e.g., One4all and MSDP) to verify the effectiveness. The result illustrates that these baselines gain huge improvement and outperform the original performance 1.1%/2.9% at AUC/KS for One4all and 0.7%/2.2% at AUC/KS for MSDP.
5.3. Ablation Study
5.3.1. Data Source
In this subsection, we conduct an ablation study to evaluate the contribution of each heterogeneous data source. The experiment involves systematically removing one data domain at a time and then pre-training a corresponding tokenizer for evaluation on our benchmarks. The results, summarized in Table 3, lead to the following key conclusions:
| Design | Volume | AUC | KS |
| 4share | 256128 | 0.8121 | 0.5007 |
| 2share+2specific | 256128 | 0.8205 | 0.5167 |
| 4share+2specific | 256128 | 0.8220 | 0.5204 |
| 3share+2specific | 1024256 | 0.8289 | 0.5332 |
5.3.2. Codebook Design
In this section, we present an ablation study to evaluate the effect of codebook architecture. Specifically, we compare a vanilla RQ-VAE with our proposed Multi-view RQ-VAE, and we additionally examine how performance scales with codebook capacity (number of layers, entries per layer, and code dimension). As shown in Table 5, the results yield two main observations:
i) Our Multi-view RQ-VAE consistently outperforms the vanilla RQ-VAE across all benchmarks. In the first two rows of Table 5 (which hold the configuration fixed at 256 entries per layer and 128-dimensional code vectors), the proposed architecture achieves superior results, validating the effectiveness of its multi-view design.
ii) Performance scales positively with increased codebook capacity. Comparing the 256x128 configurations, moving from ”2share+2specific” to ”3share+2specific” yields gains on both metrics, indicating that greater depth helps capture more complex patterns. Furthermore, increasing entries per layer from 256 to 1024 and the code dimension from 128 to 256—for example, moving from ”2share+2specific” (256x128) to ”3share+
2specific” (1024x256)—produces substantial additional improvements.
| Model | Data Volume | Training Time | Storage |
| FOUND | 20,000,000 samples | about 7 hours | 55GB |
| U2QT | 20,000,000 samples | about 2 hours | 0.65GB |
5.4. Efficiency Analysis
5.5. Codebook Analysis
5.5.1. CodeBook Utilization
In this section, we examine the utilization rates of the shared and domain-specific codebooks across the two quantization layers to better understand our model’s behavior. We define the utilization rate as the percentage of unique codes used during training, which provides insight into how effectively the model leverages its representational capacity (van der Maaten and Hinton, 2008). The results are visualized in Figure 5. From this analysis, we draw two key observations:
i) Hierarchical abstraction and shared-specific disentanglement. The observed utilization patterns indicate that our hierarchical, modular design organizes knowledge across levels of abstraction. The lower utilization of the shared codebook—especially at the first layer—suggests that a small set of cross-domain patterns is captured and separated from domain-specific ones. In contrast, the consistently higher utilization of the second layer and the domain-specific codebooks reflects the model’s ability to refine and recombine fundamental patterns into diverse, fine-grained domain representations. This synergy between hierarchical abstraction and knowledge separation supports the design’s effectiveness in representing both broad and nuanced information efficiently.
ii) Domain-dependent capacity usage. Differences in utilization across domains align with the inherent properties of the data. For example, the ”Tabular” domain exhibits the lowest utilization, which is expected because tabular data, composed primarily of numerical fields, tends to have less semantic richness and diversity than textual data from domains such as ”App” or ”Bill.” As a result, it can be represented well by a smaller, more specialized set of codes, and the observed utilization patterns faithfully reflect these underlying characteristics.
5.5.2. Multi-source Visualization
6. Conclusion
In this paper, we propose Unified User Tokenizer (U2QT), a novel framework for user representation. To unify the representation of heterogeneous user behaviors from various sources, we utilize Qwen3 embedding model to encode them in language space. For more efficient storage and computation, we design a multi-view RQ-VAE that compresses user representations into discrete tokens. We conduct comprehensive experiments across various downstream tasks to evaluate the effectiveness of our approach. The results consistently demonstrate that U2QT delivers strong representation capability, efficient storage and computation, and flexible cross-task generalization.
Appendix A Baselines